OpenStudy (hannahgrace4464):

At the city museum, child admission is $5.40 and adult admission is $9.40 . On Thursday, 167 tickets were sold for a total sales of $1189.80 . How many adult tickets were sold that day?

OpenStudy (hannahgrace4464):

not at all. i think i could figure it out if i just knew how to start it

OpenStudy (hannahgrace4464):

ok so 198,696 but what do i do with that number?

OpenStudy (hannahgrace4464):

oh yeah, i multiplied the wrong numbers sorry

OpenStudy (hannahgrace4464):

any guesses on an answer guys?

OpenStudy (hannahgrace4464):

@Sila1453 can you please help me out quickly?

OpenStudy (sila1453):

I think it is like this ; c+a=167, 5.4*c+9.4*a=1189.80 c=167-a 5.4*(167-a)+9.4*a=1189.80 solve for a